# Yixin Liu1 2  Lie Lu2  Jihui Jin2  Lichao Sun1  Andrea Fanelli2

### v2

@@ C.5 Generalization Across Standard Editing with Different Strength
For additive noise, both Gaussian and pink noise maintains 97.5% detection accuracy up to 0.1 standard deviations, though attribution degrades from 100% at low noise levels ($\leq$0.01) to 0% at high intensities ($\geq$0.5). Speed changes achieve 94.5-97.5% detection but, as expected, only maintain attribution at normal speed (1.0x). For resampling operations across frequencies from 16kHz to 1024kHz, detection remains strong (97.5%) with perfect attribution (100%). Echo effects across different delay times (0.1-0.5) consistently achieve 97.5% detection and 100% attribution, demonstrating the watermark’s comprehensive resilience to common audio transformations.
@@ C.6 Generalization Across Different Datasets
To study how XAttnMark generalizes across different domains and datasets, we conduct in-distribution and out-of-distribution tests on multiple audio datasets. We randomly sample 100 audio clips with 5s duration from each dataset. The user size $N$ is set to 1000.
@@
For out-of-distribution generalization (Table 10), we evaluate on four unseen datasets: AudioMarkBench (speech subset from Common Voice (Ardila et al., 2020)) (Liu et al., 2024a), ASVspoof (speech) (Liu et al., 2023b), MusicGen (AI-generated music) (Copet et al., 2023), and MUSDB18 (music) (Rafii et al., 2017). Our method maintains robust performance on these challenging OOD datasets, with detection accuracy ranging from 93-99% and attribution accuracy from 86-94%. The speech datasets (AudioMarkBench and ASVspoof) show slightly lower but still strong detection rates of 93-97%, while music datasets (MusicGen and MUSDB18) achieve near-perfect detection at 98-99%. Attribution performance remains consistently high across all OOD datasets, except for speed changes, which consistently fail as expected. Notably, the average attribution accuracy for OOD datasets ranges from 86-94%, demonstrating the robustness of our watermarking approach across diverse audio domains.
@@
C.6.1 Performance across the different audio duration and watermark strength $\alpha$
@@
For duration analysis, we observe that the detection accuracy remains robust (98.6–99.3%) across all tested durations from 1–10s, demonstrating our method’s effectiveness even with short audio clips. Attribution accuracy shows greater sensitivity, improving from 81.2% at 1s to a peak of 93.0% at 5s, then stabilizing at 91.5–92.6% for longer durations. This suggests 5s segments provide the optimal trade-off between practical usability and attribution performance, though reliable detection ($>$$98.5\%$) is achievable with clips as short as 1s.
For watermark strength analysis, we observe three distinct operational regimes: i) weak watermarks ($\alpha\leq 0.3$), ii) moderate watermarks ($0.4\leq\alpha\leq 0.7$), and iii) strong watermarks ($\alpha\geq 0.8$); iii) Strong watermarks ($\alpha\geq 0.8$) achieves asymptotic performance with detection $>$98.3% and attribution $>$92.3%, showing diminishing returns beyond $\alpha=1.0$.
@@
C.7 Watermark Residual Visualization
@@
Figure 9: Visualization of the original audio and watermark residuals across four different methods for a randomly selected sample from MusicCaps dataset, showing waveform (left), spectrogram (middle), and mel-spectrogram (right) representations.
C.8 Visualization of TF-weighted Penalty Distribution.
Comparison with TF-loudness loss in AudioSeal. AudioSeal adopts a coarse approach to implement auditory masking, using a naive loudness difference between the noise and the original audio. The main idea is that high loudness regions should mask lower loudness regions in the watermark residual. A positive penalty is assigned when the residual loudness is higher than the one of the original audio for each tile.
Specifically, the waveform is directly divided into $B\times W$ regions, where each region shares the same weighting term. Here, $B$ represents the number of frequency bands for splitting, and $W$ denotes the number of time tiles. The loudness difference between the watermark residual and the original waveform is then computed as $l^{w}_{b}=\text{Loudness}({{\delta}}^{w}_{b})-\text{Loudness}(x^{w}_{b})$, where $w$ is the window index and $b$ is the frequency band index. The weighting of the penalty term for each tile is computed in a per-band soft-max manner, $w^{w}_{b}=\frac{\exp(l^{w}_{b})}{\sum_{b=1}^{B}\exp(l^{w}_{b})}$. The final loundness loss is the weighted sum over all the tiles, $\ell_{Loud}=\sum_{w=1}^{W}\sum_{b=1}^{B}w^{w}_{b}\ell^{w}_{b}$.
However, we identify two limitations of this TF-based approach: (1) it lacks consideration of sophisticated auditory masking effects, where the masker and the maskee are not in the same tile; (2) using the loudness difference as the discrepancy measure only provides weak supervision for TF-guided watermarking in the spectrogram domain. To resolve these two issues, we propose to use a more sophisticated TF-weighted $\ell_{2}$ loss to guide the watermarking process. First, we simulate in the TF domain the 2D energy asymmetric decay. This helps us identifying the potential masker-maskee pairs using filtering rules based on psychoacoustic principles. Secondly, instead of using loudness difference as a discrepancy measure, we compute the mean-square error between the watermarked audio and the original one in the mel-spectrogram domain. This approach provides a more dense and fine-grained guidance signal for the imperceptible watermark.
Figure 10: Visualization of the per-tile masking penalty distribution. The blue regions indicate tiles assigned a lower penalty for embedding a watermark. Our weighting scheme provides more fine-grained control. The audio sample is randomly selected from MusicCap (Agostinelli et al., 2023), and the watermarked audio is generated from a model at the very beginning of training.

### v3

@@ C.5 Generalization Across Standard Editing with Different Strengths
For additive noise, both Gaussian and pink noise maintain 97.5% detection accuracy up to 0.1 standard deviations, though attribution degrades from 100% at low noise levels ($\leq$0.01) to 0% at high intensities ($\geq$0.5). Speed changes achieve 94.5-97.5% detection but, as expected, only maintain attribution at normal speed (1.0x). For resampling operations across frequencies from 16kHz to 1024kHz, detection remains strong (97.5%) with perfect attribution (100%). Echo effects across different delay times (0.1-0.5) consistently achieve 97.5% detection and 100% attribution, demonstrating the watermark’s comprehensive resilience to common audio transformations.
@@ C.6 Generalization Across Different Datasets
To study how XAttnMark generalizes across different domains and datasets, we conduct in-distribution and out-of-distribution tests on multiple audio datasets. We randomly sample 100 audio clips with a 5s duration from each dataset. The user size $N$ is set to 1000.
@@
For out-of-distribution generalization (Table 11), we evaluate on four unseen datasets: AudioMarkBench (Liu et al., 2024b) (which is the speech subset from Common Voice (Ardila et al., 2020)), ASVspoof (speech) (Liu et al., 2023b), MusicGen (AI-generated music) (Copet et al., 2024), and MUSDB18 (music) (Rafii et al., 2017). Our method maintains robust performance on these challenging OOD datasets, with detection accuracy ranging from 93-99% and attribution accuracy from 86-94%. The speech datasets (AudioMarkBench and ASVspoof) show slightly lower but still strong detection rates of 93-97%, while music datasets (MusicGen and MUSDB18) achieve near-perfect detection at 98-99%. Attribution performance remains consistently high across all OOD datasets, except for speed changes, which consistently fail as expected. Notably, the average attribution accuracy for OOD datasets ranges from 86-94%, demonstrating the robustness of our watermarking approach across diverse audio domains.
@@
C.7 Performance across the different audio duration and watermark strength $\alpha$
@@
For duration analysis, we observe that the detection accuracy remains robust (98.6–99.3%) across all tested durations from 1–10s, demonstrating our method’s effectiveness even with short audio clips. Attribution accuracy shows greater sensitivity, improving from 81.2% at 1s to a peak of 93.0% at 5s, then stabilizing at 91.5–92.6% for longer durations. This suggests 5s segments provide the optimal trade-off between practical usability and attribution performance, though reliable detection ($98.5\%$) is achievable with clips as short as 1s.
For watermark strength analysis, we observe three distinct operational regimes: i) weak watermarks ($\alpha\leq 0.3$), ii) moderate watermarks ($0.4\leq\alpha\leq 0.7$), and iii) strong watermarks ($\alpha\geq 0.8$); achieve asymptotic performance with detection >>98.3% and attribution >>92.3%, showing diminishing returns beyond $\alpha=1.0$.
@@
C.8 Watermark Residual Visualization
@@
Figure 11: Visualization of the original audio and watermark residuals across four different methods for a randomly selected sample from the MusicCaps dataset, showing waveform (left), spectrogram (middle), and mel-spectrogram (right) representations.
C.9 Visualization of TF-weighted Penalty Distribution.
Comparison with TF-loudness loss in AudioSeal. AudioSeal (San Roman et al., 2024) adopts a coarse approach to implement auditory masking, using a naive loudness difference between the noise and the original audio. The main idea is that high loudness regions should mask lower loudness regions in the watermark residual. A positive penalty is assigned when the residual loudness is higher than that of the original audio for each tile.
Specifically, the waveform is directly divided into $B\times W$ regions, where each region shares the same weighting term. Here, $B$ represents the number of frequency bands for splitting, and $W$ denotes the number of time tiles. The loudness difference between the watermark residual and the original waveform is then computed as $l^{w}_{b}=\text{Loudness}({{\delta}}^{w}_{b})-\text{Loudness}(x^{w}_{b})$, where $w$ is the window index and $b$ is the frequency band index. The weighting of the penalty term for each tile is computed via a soft-max over the entire time-frequency plane (San Roman et al., 2024), $w^{w}_{b}=\frac{\exp(l^{w}_{b}/\tau)}{\sum_{w^{\prime}=1}^{W}\sum_{b^{\prime}=1}^{B}\exp(l^{w^{\prime}}_{b^{\prime}}/\tau)}$, where $\tau$ is a temperature hyperparameter. The final loudness loss is the weighted sum over all the tiles, $\ell_{Loud}=\sum_{w=1}^{W}\sum_{b=1}^{B}w^{w}_{b}\ell^{w}_{b}$.
However, we identify two limitations of this TF-based approach: (1) it lacks consideration of sophisticated auditory masking effects, where the masker and the maskee are not in the same tile; (2) using the loudness difference as the discrepancy measure only provides weak supervision for TF-guided watermarking in the spectrogram domain. To resolve these two issues, we propose to use a more sophisticated TF-weighted $\ell_{2}$ loss to guide the watermarking process. First, we simulate in the TF domain the 2D energy asymmetric decay. This helps us identifying the potential masker-maskee pairs using filtering rules based on psychoacoustic principles. Secondly, instead of using loudness difference as a discrepancy measure, we compute the mean-square error between the watermarked audio and the original one in the mel-spectrogram domain. This approach provides a denser and finer-grained guidance signal for the imperceptible watermark.
Figure 12: Visualization of per-tile masking penalty distributions on the mel-spectrogram. Compared to AudioSeal, our method demonstrates more fine-grained penalty allocation. Blue colors denote regions with lower penalties, preferred for watermark embedding. The audio sample is from MusicCaps (Agostinelli et al., 2023), with the watermark generated by an early-stage model.
@@
C.10 More Results on Watermark Localization
Similar to the Brute Force Detection (BFD) in WavMark (Chen et al., 2023), we further extend our model to have the localization ability. Specifically, since our model includes transformations robust to time shifts and operates on 1s segments, we can distribute the per-segment detection probability to the per-frame level with multiple overlapping detection windows.
Mathematically, let the audio signal be represented as a sequence of frames $F=\{f_{1},f_{2},\ldots,f_{N}\}$. Our localization approach utilizes multiple overlapping detection windows, where each window $W_{k}$ corresponds to a 1s audio segment. For each window $W_{k}$, our model computes a watermark detection probability $P(M|W_{k})$, where $M$ denotes the presence of the watermark. A given frame $f_{i}$ may be part of several such overlapping windows; let this set be denoted as $\mathcal{W}(f_{i})=\{W_{k}\mid f_{i}\in W_{k}\}$. The per-frame detection probability $P(M|f_{i})$ is then determined by aggregating the probabilities from all windows covering that frame. We take the average of $P(M|W_{k})$ for all $W_{k}\in\mathcal{W}(f_{i})$ as the per-frame detection probability $P(M|f_{i})$, i.e., $P(M|f_{i})=\frac{1}{|\mathcal{W}(f_{i})|}\sum_{W_{k}\in\mathcal{W}(f_{i})}P(M|W_{k})$. This process effectively distributes the segment-level detection scores to a finer, per-frame resolution, enabling precise localization of the watermarked regions.
The results are shown in Figure 13. The experiments involve two setups: (1) fixing the audio length to 10s and varying the watermark segment length (2-10s) within it; (2) fixing the watermark segment length to 1s and varying the total audio length (2-10s). For both setups, per-frame level detection is performed. Results show that XAttnMark can achieve comparable localization performance to AudioSeal and significantly outperforms WavMark.
(a) Fixed audio length (10s), varied watermark segment length (2-10s).
(b) Fixed watermark segment length (1s), varied audio length (2-10s).
Figure 13: Watermark localization performance comparison. XAttnMark achieves localization comparable to AudioSeal and significantly outperforms WavMark.
C.11 Sensitivity Test in Model Architecture
We investigate the sensitivity of XAttnMark to key architectural choices, specifically the hidden dimension ($H$) of the message embedding table and the architecture of the temporal conditioning module. The results of this analysis, focusing on convergence speed for message decoding (message-bit validation accuracy), are presented in Figure 14.
For the message embedding table, we explored various hidden dimensions $H$ relative to the secret message bit-length $b$ (fixed at $b=16$ in our experiments). We performed a grid search over $H\in\{b/2,b,2b,4b,8b\}$. As illustrated in Figure 14(a), the choice of $H$ affects the convergence speed of the model, particularly for message decoding. We found that $H$ values ranging from $b/2$ (i.e., 8) to $4b$ (i.e., 64) yield relatively fast convergence. However, using a much larger dimension, such as $H=8b$ (i.e., 128), can slow down the convergence process (the $H=128$ case did not converge within 50k training steps).
Regarding the temporal conditioning module, we conducted an ablation study by comparing different MLP architectures: a linear projection (as proposed in XAttnMark), a 2-layer MLP, and a 3-layer MLP. The results, shown in Figure 14(b), indicate that the linear projection is the most effective, being the only one among the tested options that converges successfully within the observed training steps. This suggests that the model’s convergence is sensitive to the architectural complexity of the temporal conditioning module.
(a) Message-bit validation accuracy with different embedding dimensions ($H$).
(b) Message-bit validation accuracy with different temporal conditioning architectures.
Figure 14: Sensitivity analysis of two model architecture components within 50k training steps. The message-bit validation accuracy is evaluated on the validation set.
C.12 Inference Efficiency and Model Size
We additionally report the model size, computational complexity, training speed, and inference speed comparison with AudioSeal in Table 12. Our model uses fewer parameters and has a smaller size for both the generator and the detector. Although our generator has higher FLOPs and a slightly increased inference time per segment ($\sim$0.3 ms/segment), our detector significantly reduces FLOPs while maintaining similar inference efficiency overall. In training speed, our model achieves a similar second-per-iteration rate (around 1.15 s/iter) as AudioSeal, with faster convergence speed in learning message decoding shown in the App. C.2.
Table 12: Comparison of model statistics and inference speed between our method and AudioSeal. Training speed is computed with a batch size of 16 over the first 100 steps.
Method Component #Params $\downarrow$ Model Size $\downarrow$ FLOPs $\downarrow$ Inference Time (ms) $\downarrow$ Train Speed (s/iter) $\downarrow$
Ours Generator 10.14M 38.67 MB $1.18\times 10^{11}$ $0.411\pm 0.108$ -
Detector 7.59M 28.95 MB $1.96\times 10^{11}$ $0.981\pm 1.57$ -
Total 17.73M 67.62 MB $3.14\times 10^{11}$ $39\pm 1.58 1.1$$15\pm 0.51$
AudioSeal Generator 14.68M 56.00 MB $1.35\times 10^{10}$ $0.707\pm 0.118$ -
Detector 8.65M 32.99 MB $6.78\times 10^{11}$ $0.355\pm 0.478$ -
Total 23.33M 88.99 MB $6.92\times 10^{11}$ $1.06\pm 0.493$ $1.05\pm 0.27$
C.13 Comparison of Perceptual Quality under Different Watermark Strengths
To compare the perceptual quality of XAttnMark and AudioSeal under different watermark strengths, we conducted experiments adjusting the watermark intensity. For both XAttnMark and the AudioSeal baseline, we targeted Signal-to-Noise Ratios (SNRs) at levels from 10 dB to 50 dB, with an interval of 5 dB. The experimental results highlight two key advantages of our approach. First, as illustrated in Figure 15(a), the average attribution accuracy (across edits) of XAttnMark is consistently higher than AudioSeal when compared under different PESQ scores. Second, across different SNR levels, XAttnMark consistently better or comparable perceptual quality than AudioSeal in terms of PESQ scores. In summary, under different watermark strengths, XAttnMark achieves better/comparable performance in both attribution accuracy and perceptual quality, compared to AudioSeal.
(a) Average attribution accuracy across different PESQ constraints.
(b) PESQ scores under different SNR constraints.
Figure 15: Performance comparison under different perceptual quality constraints with targeted SNR (10-50 dB, 5 dB interval). Figure 15(a) shows that our method achieves better attribution accuracy across quality levels. Figure 15(b) shows PESQ under SNR constraints.
C.14 Statistical Significance Analysis Across Transformations
To rigorously evaluate the performance differences in robustness between XAttnMark and AudioSeal across various audio transformations, we conduct statistical significance tests. We employ McNemar’s test to compare the proportions of correct attributions (a binary outcome for paired samples) and the Wilcoxon signed-rank test to compare the distributions of bit-wise message accuracy scores (a continuous measure for paired samples). In Table 13, we report the results of these tests for each transformation applied to 10k audio samples per transformation. For the vast majority of transformations, XAttnMark demonstrates statistically significant improvements over AudioSeal. Specifically, the p-values for both McNemar’s test (on attribution correctness) and the Wilcoxon signed-rank test (on bit-wise accuracy) are typically far below the standard significance threshold of 0.05 (often $p<0.001$, indicated as 0.000 in the table). This indicates that the superior performance of XAttnMark is not a result of random variation. For instance, in the comprehensive ’Mixed of All’ scenario (consisting of all transformations), XAttnMark achieves an attribution accuracy of 0.88 compared to AudioSeal’s 0.58, and a bit-wise accuracy of 0.99 versus 0.93 for AudioSeal. In summary, XAttnMark achieves statistically significant improvements over AudioSeal across various audio transformations.
Table 13: Statistical comparison of watermarking robustness between XAttnMark and AudioSeal on our main setup (user size of $10^{4}$ is used). For each clip, we generate 100 messages and thus obtain 10k audio samples in total for each transformation. Results show both McNemar’s test on attribution correctness (discrete values) and the Wilcoxon signed-rank test on bit-wise message accuracy (continuous values). Significance is marked as ‘Yes’ for p-values << 0.05 (i.e., $\alpha=0.05$).
Transformation McNemar’s Test Wilcoxon Test
Ours AudioSeal Significant Ours AudioSeal Significant
[…]